OpenStudy (anonymous):

find the slop and y-intercept of x+y=5

OpenStudy (anonymous):

x+y=5 change them to y=mx+b (slope-intercept form) y=-x+5 then from there find the slope and the y-intercept (remember m is the slope and b is the y-intercept) imagine there's a 1 beside x so the slope is -1 and the y-intercept is 5